# Watch and Learn: Learning to Use Computers from Online Videos

### v1

@@ 4.1.2 Datasets
SFT; Labeling: TongUI 5.4 (+3.5)
SFT; IDM: W&L 13.0 (+11.1)
UI-TARS-7B (Qin et al., 2025) Base (No SFT) 27.3
SFT; Labeling: TongUI 23.8 (-3.5)
SFT; IDM: W&L 31.1 (+3.8)
Table 2: Main results on OSWorld. W&L improves general multimodal models, an agentic framework, and open-source CUAs across both in-context learning and supervised fine-tuning.
4.2 Results and Analysis
Table 2 summarizes our main results on OSWorld across both in-context learning and supervised fine-tuning. We observe consistent improvements across all model categories. For general-purpose multimodal models (Gemini 2.5 Flash, OpenAI o3, Claude 4 Sonnet), adding our W&L exemplars improves performance by +1.6 to +3.0 points. This shows that trajectories distilled from web tutorials provide useful domain-specific priors that even strong foundation models can leverage at inference time. For the Jedi agentic framework, which couples the o3 planner with Jedi grounding, W&L yields a +2.2 point gain. This demonstrates that our trajectories can complement structured planning pipelines by enriching them with exemplars that support both planning and grounding. For open-source CUAs, supervised fine-tuning on our 53k video-derived trajectories yields even larger gains. UI-TARS-7B improves by +3.8 points, while Qwen 2.5-VL sees the largest improvement, from 1.9 to 13.0 (+11.1). This larger jump is expected because Qwen is a general-purpose multimodal model not originally trained for computer use, so it benefits disproportionately from our dataset, which provides task-specific supervision that was previously missing. Overall, these results highlight the value of our dataset as a scalable supervision signal for both specialized CUAs and broader multimodal models.
@@ 4.2 Results and Analysis
Figure 3: Qualitative examples on OSWorld. On the left, the video-derived trajectory that W&L generates for the task. On the right: (i) the o3 agent makes a grounding error by selecting a wrong UI element; (ii) the Jedi (o3) agent makes a planning error by entering the wrong submenu without recovering; (iii) using the video-derived trajectory, W&L agent completes the task successfully. Images are cropped for visibility, and the action coordinates correspond to the original full-resolution screenshots.
@@
Table 3: Ablation study on the effect of action labeling and reasoning in ICL exemplars (OSWorld success rates). Structured trajectories provide consistent gains over raw frames across all models.
ActionType Gemini 2.5 Flash TongUI W&L IDM
click(x, y) 69.2% 72.7% 94.4%
scroll(scroll_y) 70.5% 76.4% 93.7%
type(text) 77.2% 71.8% 78.5%
wait(500ms) 92.3% 94.1% 97.5%
move(x, y) 65.8% 70.3% 89.2%
Action Accuracy 72.8% 82.7% 91.6%
ActionType Accuracy 81.4% 88.9% 96.4%
Table 4: Comparison of action labeling accuracy on the Mind2Web test set. W&L’s IDM outperforms TongUI, achieving the best performance
4.2.1 How much do labeled trajectories help in in-context learning?
We next analyze the contribution of accurate video labeling to in-context learning (ICL). Our framework provides structured action annotations and natural language reasoning for each step. To isolate the effect of each, we compare three variants: (i) consecutive frames only, (ii) frames paired with predicted actions, and (iii) frames with both actions and reasoning generated by Gemini 2.5 Flash.
Ablations on OSWorld (Table 3) show that adding action labels provides a substantial boost over using frames alone, and further gains are achieved when natural language reasoning is included. This pattern holds consistently across all tested models. Figure 3 provides a qualitative example, showing how labeled trajectories impact the original agent’s behavior. The improvement demonstrates that labeled trajectories do more than supply visual context; they encode procedural knowledge that helps models improve both planning and grounding for complex workflows.
4.2.2 How does label accuracy impact performance?
Action label accuracy is central to training CUAs: noisy annotations not only fail to help but can actively degrade performance. We first compare our dedicated IDM against Gemini 2.5 Flash and the TongUI labeling pipeline (based on UI-TARS-7B) on the held-out Mind2Web test set (Table 4).
Our IDM achieves the strongest results, substantially outperforming both baselines. TongUI offers some gains over Gemini, especially for structured actions such as scroll and click, but still falls short of our IDM. A remaining limitation is text decoding for type actions, where the margin is smaller.
These differences in labeling accuracy directly translate into downstream performance. TongUI, despite sharing our prompt format, relies on noisy labels that hurt both in-context learning and fine-tuning (Table 4). With o3, TongUI exemplars reduce success rates; in model training, they yield only marginal gains for Qwen and even lower UI-TARS performance (Table 2). In contrast, our IDM-derived labels consistently improve performance, underscoring that reliable supervision is key for effective action grounding.
4.2.3 What is the effect of retrieval quality for in-context learning?
@@
Table 5: ICL results on OSWorld with o3. Random retrieval has little effect, while W&L yields strong gains.
@@ 5 Conclusion and Future Work
We introduced W&L, a framework that transforms web-scale human demonstration videos into executable UI trajectories using a vision-only IDM and a task-aware retrieval pipeline. With over 53k automatically labeled trajectories, we showed improvements in both in-context learning and supervised fine-tuning, benefiting general-purpose MLLMs as well as specialized CUAs.
Our experiments on OSWorld highlight that (i) a dedicated IDM provides stronger action prediction than foundation models, (ii) action-labeled exemplars improve ICL and SFT, (iii) domains with abundant tutorials see larger gains, and (iv) performance scales with more training data and better retrieval.
Looking ahead, we plan to extend the IDM to richer actions such as drag-and-drop, combine or split tutorials to better construct long-horizon trajectories, and explore reinforcement learning with our trajectories—using them as demonstrations for behavior cloning, as replay buffers for offline RL, or as priors for reward modeling in online training. These directions can further bridge large-scale demonstrations with adaptive learning, pushing CUAs closer to real-world deployment.
@@ Acknowledgments
Appendix A Use of Large Language Models (LLMs)
Large language models (LLMs) were used in limited ways during this work. Specifically, we used LLM-based assistants to (i) improve sentence structure, paragraph organization, and grammar in the writing process, and (ii) provide coding assistance such as debugging and suggesting alternative implementations. LLMs were not used for research ideation, experimental design, or analysis. All scientific contributions, including problem formulation, methodology, experiments, and conclusions, are solely the work of the authors.
Appendix B Limitations
While our framework demonstrates strong performance, there remain several opportunities for extension. First, our inverse dynamics model (IDM) currently focuses on a core set of primitive actions such as click, type, move, and scroll. More complex actions like drag-and-drop are not yet supported, largely due to the absence of sufficient training data. Similarly, while our IDM predicts scroll actions, we were unable to curate a large-scale, diverse dataset of scrolling behaviors from web interactions, which limits its robustness in this dimension. Expanding the action space with richer interaction types and collecting more representative scroll data are promising directions.
@@ Appendix B Limitations
Appendix C Dataset Details
C.1 Applications by Category
@@
Design Adobe Photoshop, Adobe Illustrator, Adobe XD, Figma, Sketch, Canva, CorelDRAW, Inkscape, Affinity Designer
@@
Table 6: Applications grouped by category.
Appendix D Implementation Details
D.1 Video Retrieval
@@ D.2 Video Filtering
D.3 Models
@@
Appendix E More Results
E.1 What is the effect of data scale for supervised fine-tuning?
Model Base 10k 25k Full
Qwen 2.5-VL 1.9 3.3 4.9 13.0
E.2 Which application domains benefit most from our data?

### v2

@@ 3.2 Trajectory Data Generation from Videos
SFT w/ TongUI labeled 5.4 (+3.5)
SFT w/ W&L (IDM labeled) 13.0 (+11.1)
UI-TARS-1.5-7B (Qin et al., 2025) Base (No SFT) 27.3
SFT w/ TongUI labeled 23.8 (-3.5)
SFT w/ W&L (IDM labeled) 31.1 (+3.8)
Table 3: WindowsAgentArena results under the 15-step evaluation limit. Comparison of open-source 7B-sized CUAs. W&L achieves the highest success rate among 7B-sized models, surpassing both prior automatic labeling methods and zero-shot baselines.
Model Setting / Training Data Success Rate (%)
UI-TARS-1.5-7B (Qin et al., 2025) Zero-shot (no SFT) 18.1
SFT w/ TongUI (Zhang et al., 2025) 12.9 (-5.2)
SFT w/ W&L (IDM-labeled) 24.0 (+5.9)
OpenCUA-7B (Wang et al., 2025) Base 13.5
UltraCUA-7B (Yang et al., 2025b) Base 21.7
3.3 Applications of Trajectories
The extracted trajectories are utilized in two settings. For in-context learning (ICL), each trajectory is converted into a set of (observation, action) pairs augmented with concise natural-language reasoning traces generated by Gemini 2.5 Flash . These (observation, action, reasoning) exemplars are formatted as few-shot demonstrations in the model context, enabling general-purpose agents to leverage planning and grounding priors without additional training. For supervised fine-tuning (SFT), the labeled trajectories are aggregated into a large corpus of (state, action) sequences and used to optimize multimodal LLMs with a standard sequence modeling objective. We fine-tune UI-TARS-1.5 (Qin et al., 2025), a specialized computer-using agent, and Qwen 2.5-VL (Bai et al., 2025), a general multimodal model, demonstrating that the proposed data serve as a scalable and transferable supervision source across diverse models.
@@ 3.3 Applications of Trajectories
Figure 3: Qualitative examples on OSWorld. On the left, the video-derived trajectory that W&L generates for the task. On the right: (i) the o3 agent makes a grounding error by selecting a wrong UI element; (ii) the Jedi (o3) agent makes a planning error by entering the wrong submenu without recovering; (iii) using the video-derived trajectory, W&L agent completes the task successfully. Images are cropped for visibility, and the action coordinates correspond to the original full-resolution screenshots. More trajectory examples are in Appendix.
4 Experiments and Analysis
4.1 Experimental Setup
Our experiments evaluate whether IDM-labeled trajectories improve interactive computer-using agents in two settings: (1) in-context learning (ICL) for general-purpose multimodal models and agentic frameworks, and (2) supervised fine-tuning (SFT) for open-source models. We also compare the resulting trajectories against TongUI (Zhang et al., 2025), an existing automatic labeling pipeline.
All agents operate under a unified interface where each step provides only the current screen image and the history of executed actions. In the ICL setting, each model receives a fixed number of task-relevant trajectories retrieved from our video-derived dataset as exemplars. In the SFT setting, open-source CUAs are trained directly on the same IDM-labeled trajectories.
We evaluate on two complementary interactive computer-use benchmarks. OSWorld-Verified (Xie et al., 2024) provides full desktop and operating-system (Ubuntu) environments across productivity, programming, design, and system utilities, and we apply a 50-step interaction limit. To assess robustness across different interface conventions and OS-specific workflows, we additionally evaluate on WindowsAgentArena (Bonatti et al., 2024), a benchmark of Windows tasks that follows a 15-step interaction limit. Together, these benchmarks allow us to test whether the benefits of IDM-labeled trajectories generalize across heterogeneous operating systems and UI distributions.
4.2 Datasets
State-transition corpus. To train the IDM, we gather autonomous state transitions from real web interactions, resulting in over 600k $(O_{t},a_{t},O_{t+1})$ triples.
Video-derived trajectories. After training the IDM, we apply it to a curated collection of 53,125 instructional YouTube videos, producing one trajectory per video and yielding 53,125 trajectories across 69 applications spanning productivity, programming, design, audio production, system utilities, and scientific/data domains (Table 1). Additional details are provided in the Appendix.
Baseline labeler. TongUI (Zhang et al., 2025) generates action annotations by prompting the UI-TARS-1.5-7B agent (Qin et al., 2025), while our method applies the trained IDM to the same source videos. To ensure a fair comparison focused only on labeling quality, we use the identical set of tutorial videos retrieved by our framework and apply TongUI’s labeling pipeline, differing only in the model used to annotating actions.
4.3 Models
We evaluate three categories of models covering general multimodal model, agentic framework, and open-source CUAs.
General-purpose multimodal models. Gemini 2.5 Flash (Gemini Team, 2025), OpenAI o3 (OpenAI, 2025a), and Claude 4 Sonnet (Anthropic, 2025) are evaluated in the ICL setting using retrieved IDM-labeled trajectories as exemplars.
Agentic framework. We use Jedi (Xie et al., 2025), a vision-only agentic framework for OSWorld that combines an MLLM planner (OpenAI o3) with the Jedi-7B grounding model that maps natural language steps to executable UI actions. Its modular design allows straightforward integration of our ICL exemplars, so we evaluate Jedi both with and without retrieved IDM-labeled trajectories.
Open-source CUAs. UI-TARS-1.5-7B (Qin et al., 2025) and Qwen 2.5-VL 7B (Bai et al., 2025) are trained with supervised fine-tuning on our 53,125 video-derived trajectories, enabling a controlled comparison against TongUI-labeled data and the exemplar-based performance of multimodal models.
4.4 Main Results
Table 2 summarizes our main results on OSWorld across both in-context learning and supervised fine-tuning. We observe consistent improvements across all model categories. For general-purpose multimodal models (Gemini 2.5 Flash, OpenAI o3, Claude 4 Sonnet), adding our W&L exemplars improves performance by +1.6 to +3.0 points. This shows that trajectories distilled from web tutorials provide useful domain-specific priors that even strong foundation models can leverage at inference time. For the Jedi agentic framework (Xie et al., 2025), which couples the o3 planner with Jedi-7B grounding model, W&L yields a +2.2 point gain, demonstrating that our trajectories can complement structured planning pipelines by enriching them with exemplars that support both planning and grounding. For open-source CUAs, supervised fine-tuning on our 53k video-derived trajectories yields even larger gains. UI-TARS-7B improves by +3.8 points, while Qwen 2.5-VL sees the largest improvement, from 1.9 to 13.0 (+11.1). This larger jump is expected because Qwen is a general-purpose multimodal model not originally trained for computer use, so it benefits disproportionately from our dataset, which provides task-specific supervision that was previously missing.
We further evaluate on WindowsAgentArena (Bonatti et al., 2024), which focuses on Windows OS tasks under a standardized 15-step evaluation limit. As shown in Table 3, W&L achieves the state-of-the-art success rate among 7B-scale models with 15 step limit, surpassing both prior automatic labeling method and specialized CUA baselines. Specifically, fine-tuning UI-TARS-1.5-7B on IDM-labeled trajectories improves performance from 18.1 to 24.0 (+5.9), whereas training on TongUI-labeled data leads to a degradation in performance (12.9, –5.2). These results confirm that IDM-based labeling produces more accurate and consistent supervision, translating to stronger downstream generalization.
Overall, across both OSWorld and WindowsAgentArena, W&L serves as a scalable and high-quality supervision signal that consistently enhances models and agentic frameworks in both in-context and fine-tuning settings.
4.4.1 Effect of Label Accuracy on Performance
Action label accuracy is central to training CUAs: noisy annotations not only fail to help but can actively degrade performance. We first compare our dedicated IDM against Gemini 2.5 Flash and the TongUI labeling pipeline (based on UI-TARS-1.5-7B) on the held-out state-transition test set (Table 4). Our IDM achieves the strongest results, substantially outperforming both baselines. TongUI offers modest gains over Gemini, especially for structured actions such as scroll and click, but still falls short of our IDM.
These differences in labeling accuracy directly translate into downstream performance. TongUI, although it follows the same prompt format, relies on labels generated by UI-TARS-1.5-7B, whose own performance is notably weaker on WindowsAgentArena (Table 3) than on Ubuntu-based OSWorld (Table 2). As a result, labeling quality degrades more severely in the Windows setting, leading to noisier supervision. As shown in Table 3, this effect is amplified on WindowsAgentArena, where fine-tuning with TongUI-labeled data causes a drop in success rate (a decrease of 5.2 points) compared with consistent gains from IDM-labeled trajectories. In contrast, our IDM-derived labels remain reliable across both Ubuntu-style (OSWorld) and Windows-style (WindowsAgentArena) interfaces, underscoring that accurate and platform-robust supervision is essential for effective action grounding and cross-OS generalization.
4.4.2 Effect of Trajectories on In-Context Learning
We next analyze the contribution of accurate video labeling to in-context learning (ICL). Our framework provides structured action annotations and natural language reasoning for each step. To isolate the effect of each component, we compare three variants: (i) consecutive frames only, (ii) frames paired with predicted actions, and (iii) frames with both actions and reasoning generated by Gemini 2.5 Flash.
Ablations on OSWorld (Table 5) show that adding action labels provides a substantial boost over using frames alone, and further gains are achieved when natural language reasoning is included. This pattern holds consistently across all tested models. Figure 3 illustrates these effects qualitatively: baseline agents without labeled trajectories often make grounding or planning errors, for example clicking incorrect UI elements or entering the wrong submenu, while the W&L-enhanced agent successfully completes the same tasks by leveraging accurate video-derived exemplars. These observations confirm that labeled trajectories convey procedural and causal knowledge beyond visual context, improving both planning and grounding for complex workflows.
Table 4: Action labeling accuracy on the held-out test set. We evaluate six primitive action types with 100 transitions each. ActionType Accuracy reflects correct action classification, while Action Accuracy also checks argument correctness. W&L’s IDM outperforms Gemini and TongUI across all action types.
ActionType Gemini 2.5 Flash TongUI W&L IDM
click(x, y) 68% 72% 95%
release 71% 67% 90%
scroll(scroll_y) 55% 75% 93%
type(text) 77% 71% 86%
wait(500ms) 92% 88% 97%
move(x, y) 65% 61% 89%
ActionType Accuracy 81.5% 84.3% 95.8%
Action Accuracy 70.5% 72.3% 91.7%
Table 5: Ablation of action labeling and reasoning within ICL exemplars on OSWorld. Structured trajectories lead to steady gains compared with raw frame exemplars.
@@ 4.4.2 Effect of Trajectories on In-Context Learning
4.4.3 Effect of Video Retrieval on In-Context Learning
Table 6: Ablation of retrieval strategy for ICL on OSWorld. Random retrieval gives minimal gains, while task-relevant retrieval in W&L improves performance.
@@ 5 Conclusion and Future Work
We introduced W&L, a framework that converts web-scale human demonstration videos into executable UI trajectories using a vision-only inverse dynamics model (IDM) and a task-aware video retrieval pipeline. With over 53k automatically labeled trajectories, W&L improves both in-context learning and supervised fine-tuning across diverse agents, benefiting general-purpose multimodal models as well as open-source CUAs. Experiments on OSWorld and WindowsAgentArena confirm that IDM-labeled trajectories provide high-quality supervision that generalizes across operating systems and interface conventions.
Future work will focus on combining or segmenting tutorials to better capture long-horizon workflows and on exploring reinforcement learning using our trajectories as demonstrations, replay buffers, or priors for reward modeling. These directions aim to further bridge large-scale demonstrations with adaptive policy learning, bringing CUAs closer to robust real-world deployment.
@@ Acknowledgments
Appendix A Limitations
While our framework demonstrates strong performance, several opportunities for extension remain. Our inverse dynamics model (IDM) currently focuses on a core set of primitive actions such as click, type, move, and scroll. It does not yet handle keyboard shortcut or hotkey based actions (e.g., Ctrl+C or Ctrl+Z), as such events are rarely captured in web based state transition data and therefore lack corresponding visual context. Nevertheless, most hotkey operations have direct graphical interface equivalents, such as selecting menu items or clicking toolbar icons, allowing the IDM to retain near complete functional coverage of real world tasks. Future work could incorporate application based data collection or key event logging to expand the action space and support shortcut based interactions more efficiently.
@@ Appendix A Limitations
Appendix B Dataset Details
B.1 Applications by Category
@@
Design Adobe Photoshop, Adobe Illustrator, Adobe XD, Figma, Sketch, Canva, CorelDRAW, Inkscape, Gimp
@@
Appendix C Implementation Details
C.1 State-Transition Data Collection
This section provides additional details of the automated pipeline used to gather the 600K state-transition pairs $(O_{t},a_{t},O_{t+1})$ described in the main paper. The system consists of four components: an entry-point sampler, a browser controller, an action executor, and a transition logger. All data were collected using Playwright with a fixed viewport of $1280\times 720$ pixels and 48 parallel workers.
Entry-point sampling. We sample URLs from the March 2025 snapshot of the Common Crawl index  [^2] https://commoncrawl.org/blog/march-2025-crawl-archive-now-available (CC-MAIN-2025-13). We retain only HTML documents, filter out pages smaller than 5 KB, and discard URLs that trigger login walls or CAPTCHA challenges. To reduce domain imbalance, we cap the number of visits to 30 pages per domain. Each browsing session begins from a randomly chosen entry point, ensuring high visitation diversity across the crawl.
Action sampling policy. At each timestep, the pipeline samples an action from our six primitives: click, release, scroll, type, wait, and move. We empirically bias the sampling distribution toward common web interactions: click ($\sim$45%), scroll ($\sim$20%), move ($\sim$15%), type ($\sim$8%), wait ($\sim$2%), and release ($\sim$10%). Coordinates for click and move actions are drawn uniformly over the viewport and jittered around visible UI elements obtained from Playwright’s accessibility tree. Scroll magnitudes follow a fixed value (500 px). Text for type actions is sampled from a curated list of $\sim$5K frequently used English words  [^3] https://www.vocabulary.com/lists/154147.
Transition recording. For each action, we record the pre- and post-action screenshots $(O_{t},O_{t+1})$, the action primitive, all associated parameters, the current page URL, and a timestamp. Transitions where $O_{t}$ and $O_{t+1}$ differ by fewer than 50 nonzero pixels (excluding wait) are removed to avoid degenerate interactions. Each action is executed with a 5 s timeout; failed executions or delayed page loads are logged and the transition is discarded.
Error handling and session resets. The system automatically detects HTTP 4xx/5xx responses, navigation loops, frozen page states, and browser crashes. Upon such events, the session is terminated and restarted with a fresh entry point. To promote diversity, each session is limited to 15 actions before being reset, preventing oversampling from any single page.
Scale and runtime. Data collection was run on CPU-only machines (32 cores, 64 GB RAM). Across 48 parallel workers, the pipeline collected approximately 600K high-quality transitions over 3.5 days. After filtering, the resulting dataset covers a wide range of layouts, domains, and dynamic interaction patterns encountered in real web environments.
C.2 Video Retrieval
@@ C.3 Video Filtering
C.4 In-context Learning
To evaluate W&L under in-context learning (ICL), we augment the default prompts used by state-of-the-art OSWorld agents with a small number of structured demonstration trajectories. The OSWorld repository provides model-specific system prompts in OS-World Github  [^4] https://github.com/xlang-ai/OSWorld/blob/main/mm_agents/prompts.py, which define the action format, reasoning style, and interface conventions. We prepend our exemplars to these prompts without modifying any downstream agent logic.
For a given evaluation task, we retrieve three relevant trajectories from our video corpus and convert each into a sequence of (observation, thought, action) triples. These triples are formatted using the exact conventions defined by OSWorld agents, including the same observation style, reasoning structure, and action specification. We prepend the resulting exemplars to the agent prompt by placing them inside a dedicated instruction block that appears immediately before the OSWorld rules and reasoning instructions. This placement ensures that the agent first learns the desired behavior pattern from the exemplars, and then applies the original constraints that follow.
ICL Example Prompt Block Example Trajectories Below are some example task trajectories showing how to observe, think, and act. Study these examples carefully to understand: how to provide detailed observations of the screen state, how to structure your reasoning process step by step, how to format action code with clear comments, how to choose appropriate coordinates based on visual elements, when to use special codes like WAIT, DONE, or FAIL, how to recover from errors by adjusting your approach. These examples demonstrate similar tasks to what you’$l$ be performing. Use them as a reference for the level of detail expected in your observations and the reasoning process for choosing actions. Pay attention to how coordinates are selected based on element positions and how actions are adapted when previous attempts don’t succeed. Example {example_num}: {task_description} Step {step_num}: Screenshot: [Image showing the current state] Thought: {thought_text} Action:
{action_code}
Example of Full ICL Prompt with O3 Agent. In section˜F.2 we show how a set of exemplar trajectories is prepended to the O3 agent instructions. The exemplars appear after the initial instruction, followed immediately by the rule and reasoning sections used by the OSWorld agent.
C.5 Models
@@
Appendix D Inverse Dynamics Model Architecture
D.1 Vision Encoder
Model Selection. We employ SigLIP-2 Base (patch size 16) with the NaFlex variant as our vision encoder [Tschannen et al., 2025]. This architecture was specifically chosen for its superior performance on OCR and document-understanding benchmarks, which closely aligns with the requirements of computer interface understanding where precise text recognition and spatial localization are critical.
Input Processing. Each observation $O_{t}$ is a $\times1000$ RGB screenshot. We utilize the NaFlex preprocessing strategy with max_num_patches=1024, which preserves the square aspect ratio while resizing the input to approximately $\times704$ pixels before patch extraction. This configuration yields exactly 1024 visual tokens (a $\times32$ grid of $\times16$ patches), providing sufficient spatial resolution for identifying small UI elements such as buttons, icons, and text labels while maintaining computational efficiency.
Feature Extraction. For the pair $(O_{t},O_{t+1})$, we process each observation independently through the frozen SigLIP-2 encoder to obtain visual embeddings $\mathbf{v}_{t},\mathbf{v}_{t+1}\in\mathbb{R}^{1024\times 768}$. These embeddings preserve spatial information critical for coordinate prediction.
Rationale for NaFlex. Compared to the fixed-resolution SigLIP-2 variant at $\times512$ (which would downsample our $\times1000$ input by 48%), NaFlex provides: (1) Higher effective resolution ($\sim$$\times704$) for better preservation of fine-grained UI details, (2) Superior performance on text-heavy and structured visual content [Tschannen et al., 2025], and (3) Minimal aspect ratio distortion for square inputs. While fixed-resolution variants benefit from additional distillation stages, preliminary experiments showed NaFlex’s advantages in OCR-like capabilities outweigh this trade-off for UI understanding.
D.2 Transformer Backbone
Temporal Feature Fusion. The patch-level features from $(O_{t},O_{t+1})$ are concatenated along the sequence dimension, yielding a combined representation of shape $\mathbb{R}^{2048\times 768}$. We prepend a learnable [CLS] token to this sequence. To distinguish temporal information, we add learnable temporal embeddings $\mathbf{e}_{t},\mathbf{e}_{t+1}\in\mathbb{R}^{768}$ to the respective halves of the sequence.
Architecture. Our backbone consists of 4 transformer layers with the following specifications:
  • Hidden dimension: 768 (matching SigLIP-2 output)
  • Attention heads: 12
[…]
  • Activation function: GELU
  • Dropout: 0.1 applied to attention weights and MLP outputs
  • Layer normalization: Pre-norm configuration with $\epsilon=10^{-6}$
  • Attention mechanism: Standard scaled dot-product attention with causal masking disabled (full bidirectional attention)
The final [CLS] token representation $\mathbf{h}_{\text{cls}}\in\mathbb{R}^{768}$ serves as the global action embedding, while the patch-level outputs $\mathbf{h}_{\text{patch}}\in\mathbb{R}^{2048\times 768}$ are used for coordinate prediction.
D.3 Prediction Heads
D.3.1 Action Classification Head
A simple two-layer MLP predicts the action type:
$\mathbf{p}_{\text{action}}=\text{softmax}(\mathbf{W}_{2}\cdot\text{ReLU}(\mathbf{W}_{1}\mathbf{h}_{\text{cls}}+\mathbf{b}_{1})+\mathbf{b}_{2})$ (1)
where $\mathbf{W}_{1}\in\mathbb{R}^{512\times 768}$, $\mathbf{W}_{2}\in\mathbb{R}^{6\times 512}$, and $\mathbf{p}_{\text{action}}\in\mathbb{R}^{6}$ represents the probability distribution over the six action primitives: click, release, scroll, type, wait, and move.
D.3.2 Coordinate Prediction Head
Architecture. For location-based actions, we predict normalized coordinates discretized into 1000 bins for each axis. We employ separate prediction heads for $x$ and $y$ coordinates:
For the $x$-coordinate:
$\mathbf{h}_{x}$ $=\text{AvgPool}(\mathbf{h}_{\text{patch}})\quad\text{(over spatial dimension)}$ (2)
$\mathbf{p}_{x}$ $=\text{softmax}(\mathbf{W}_{x}\cdot\text{ReLU}(\text{LayerNorm}(\mathbf{h}_{x})))$ (3)
where $\mathbf{W}_{x}\in\mathbb{R}^{1000\times 768}$ and $\mathbf{p}_{x}\in\mathbb{R}^{1000}$. The $y$-coordinate head is parameterized identically with separate weights $\mathbf{W}_{y}$.
Discretization Strategy. Screen coordinates are normalized to $[0,1]$ and discretized into 1000 uniform bins. This formulation offers several advantages: (1) Transforms regression into classification, enabling the use of cross-entropy loss which empirically provides stronger gradients than regression losses (e.g., L1, L2), (2) Naturally handles multi-modal distributions when multiple valid click locations exist, (3) Provides calibrated uncertainty estimates via the output probability distribution.
Resolution Alignment. The 1000-bin discretization aligns naturally with our $\times1000$ input resolution, where each bin corresponds to approximately a 1-pixel region. This granularity is sufficient for UI interaction, as most clickable targets span multiple pixels.
D.3.3 Language Head
Architecture. For text-entry actions (type), we employ a lightweight autoregressive decoder based on GPT-2 [Radford et al., 2019].
Visual Conditioning. The decoder attends to the visual features via cross-attention. Specifically, at each decoder layer $\ell$, we insert a cross-attention sublayer:
$\text{CrossAttn}^{(\ell)}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (4)
where $Q$ is derived from the decoder’s hidden states, and $K,V$ are projections of the concatenated visual features $\mathbf{v}_{t};\mathbf{v}_{t+1}]$ from the SigLIP-2 encoder. This design enables the model to ground its text generation in the observed visual transition.
Training Strategy. During training, we use teacher forcing with cross-entropy loss on the ground-truth text sequence.
Appendix E Inverse Dynamics Model Training
This part provides the full definition of the training losses used for the inverse dynamics model (IDM) in W&L. The IDM predicts three types of supervision depending on the action: an action primitive, optional pixel coordinates, and optional text tokens. For completeness, we describe the outputs, losses, and training objective in detail below.
E.1 Model Outputs
Given two consecutive observations $(O_{t},O_{t+1})$, where $O_{t}$ and $O_{t+1}$ are screen images before and after executing action $a_{t}$, the IDM produces the following outputs:
  • Action primitive: A categorical distribution $p_{\theta}(a_{t})$ over the set of supported primitives (e.g., click, scroll, type, wait, move).
  • Coordinate predictions: For actions requiring a spatial position, the model predicts distributions $p_{\theta}(x_{t})$ and $p_{\theta}(y_{t})$ over discretized pixel bins $\{0,\ldots,999\}$.
  • Typed text: For text-entry actions, the IDM predicts the sequence of typed tokens $w_{1:L}$ autoregressively through $p_{\theta}(w_{\ell}\mid w_{<\ell},O_{t},O_{t+1})$.